Suppose a treadmill has an average acceleration of 0.0047 m/s^2. How much does its speed change after 4.7 seconds?
1
Expert's answer
2019-09-23T09:14:13-0400

We can use law for speed in case of uniformly accelerated motion


"v = {v_0} + at"

For the change of speed we get

"\Delta v = v - {v_0} = at = 0.0047[\frac{m}{{{s^2}}}] \cdot 4.7[s] = 0.02209[\frac{m}{s}]"